OpenStudy (anonymous):

What is the inverse of y=3x+9?

OpenStudy (anonymous):

Let f ( x ) = 3x + 9 = y and g ( y ) = x ie g is inverse of f 3x + 9 = y 3x = y - 9 x = (1/3) y - 3 g ( y ) = (1/3) y - 3 g ( x ) = (1/3) x - 3 is inverse function